OpenStudy (anonymous):

Please take the time to read the whole thing!! I wrote down an equation is class, but i only wrote down the working out part. Not acually HOW to do the working out. The Question is- 5(x-7)=25 The Answer is- x=12 My working out is. 5(x-7)=25 *underneath* 5x-35=25 *underneath* 5x=60 *underneath* x=12. Please Help

OpenStudy (diyadiya):

YOu've done it correctly

OpenStudy (anonymous):

I know but i dont know HOW i did it :/

OpenStudy (diyadiya):

5x-35=25 Add 35 on both sides 5x-35+35 =25+35 5x=60 Divide booth sides by 5\[\frac{5x}{5}=\frac{60}{5}\]x=12

OpenStudy (diyadiya):

is it clear ? is this what you mean by 'how'? @Myk

OpenStudy (anonymous):

Yes i understand it all EXCEPT how you got 5x-35=25, from, 5(x-7) = 25.

OpenStudy (diyadiya):

Oh a(b+c)= ab+ac :)